MICROWAVE SPECTRUM OF HEXAFLUOROISOPROPANOL

\underline{ABHISHEK SHAHI}, DEVENDRA MANI and E. ARUNAN, Department of Inorganic and Physical Chemistry, Indian Institute of Science, Bangalore-560012, India.

Hexafluoroisopropanol (HFIP) is an important organic solvent and probably the only solvent which can dissolve polythene. IR studies, on this molecule confirm the existence of antiperiplanar (ap) and synclinical (sc) conformers^{a}. We have observed pure rotational spectrum of this molecule and the fitted rotational constants (A= 2105.1208(11) MHz, B= 1053.9942(3) MHz, C= 932.3398(3) MHz) confirm the presence of ap conformer. There are many other observed lines which most probably corresponds to sc structure and due to the large amplitude motion of H-atom, some of these transitions show tunneling splitting. Work is in progress for the deuterated (OD) and C-13 isotopologues of the monomer. HFIP is expected to exhibit interesting hydrogen bonding properties and we are planning to investigate them by studying its complex with water. The results will be presented in this talk.

aH. Schaal, T. Höber, and M. A. Suhm, J. Phys. Chem. A 104, 265 (2000).